OpenStudy (anonymous):

0.088m/s convert to km/hour

OpenStudy (nincompoop):

there are 1,000 metres in a kilometre and there are 60 seconds in a minute and 60 minutes in an hour try to work this out

OpenStudy (anonymous):

\[0.088*\frac{ 18 }{ 5 }km/h\]

OpenStudy (nincompoop):

ermm

OpenStudy (anonymous):

LOL

OpenStudy (kropot72):

There are 60 * 60 = 3600 seconds in an hour. First find out how many meters will be travelled in one hour by multiplying by 3600. Distance travelled in 1 hour = 0.088 * 3600 meters. Next convert the distance to kilometers by dividing by 1000. So the speed in kilometers per hour is given by: \[0.88\times \frac{3600}{1000}=(you\ can\ calculate)\ km/hour\]